# Unsupervised and Source-Free Ranking of Biomedical Segmentation Models

### v2

@@ 11.2 Nuclei Perturbation Sweep
11.3 Baseline Transferability Metrics
We also show the correlation scores per dataset for all the baseline transferability metrics. Tab. 19 shows the results for the four mitochondria datasets (EPFL, Hmito, Rmito and VNC) and Tab. 20 shows the results for the four nuclei datasets (BBBC039, DSB2018, Hoechst, S_BIAD895). In the main paper due to space constraints we showed the correlation scores averaged within the dataset groups.
@@ 11.3 Baseline Transferability Metrics
11.4 Multiclass Toothfairy2 Perturbation Sweep
For multiclass semantic segmentation we investigated applying both additive Gaussian noise (controlled by ) and Gamma Correction (controlled by ) as input perturbations. For feature space perturbation we investigated applying spatial DropOut to the bottleneck and skip connection layers of the network, where the strength of the perturbation is controlled by the proportion of feature maps dropped at each layer . We report the performance of both CTE-EI (Tab. 21) and CTE-NHD (Tab. 22).
@@ 11.4 Multiclass Toothfairy2 Perturbation Sweep
| CTE-EI Gauss ( 2.5x) | K | 0.90 | (**) | 0.81 | (**) | 0.62 | (0.1) | 0.43 | (0.3) | 0.71 | (*) |
| S | 0.96 | (**) | 0.89 | (*) | 0.82 | (*) | 0.50 | (0.3) | 0.86 | (*) | |
| P | 0.99 | (**) | 0.94 | (**) | 0.70 | (0.1) | 0.38 | (0.4) | 0.93 | (**) | |
| CTE-EI Gauss ( 3.0x) | K | 0.90 | (**) | 0.81 | (*) | 0.62 | (0.1) | 0.43 | (0.2) | 0.71 | (*) |
| S | 0.96 | (**) | 0.89 | (**) | 0.82 | (*) | 0.50 | (0.3) | 0.86 | (*) | |
| P | 0.99 | (**) | 0.92 | (**) | 0.63 | (0.1) | 0.39 | (0.4) | 0.92 | (**) | |
| CTE-EI Gamma ( 0.2) | K | 0.90 | (**) | 0.71 | (*) | 0.62 | (0.1) | 0.52 | (0.2) | 0.81 | (**) |
| S | 0.96 | (**) | 0.86 | (*) | 0.75 | (0.1) | 0.57 | (0.2) | 0.89 | (*) | |
| P | 0.99 | (**) | 0.68 | (0.1) | 0.61 | (0.2) | 0.46 | (0.3) | 0.73 | (0.1) | |
| CTE-EI Gamma ( 0.5) | K | 1.00 | (**) | 0.81 | (**) | 0.81 | (*) | 0.43 | (0.3) | 0.81 | (**) |
| S | 1.00 | (**) | 0.93 | (**) | 0.89 | (**) | 0.43 | (0.3) | 0.89 | (**) | |
| P | 0.96 | (**) | 0.91 | (**) | 0.81 | (*) | 0.44 | (0.3) | 0.91 | (**) | |
| CTE-EI Gamma ( 2.0) | K | 1.00 | (**) | 0.71 | (*) | 0.43 | (0.3) | 0.43 | (0.2) | 0.81 | (**) |
@@
11.5 Post-UDA Perturbation Sweep
In this section we investigate the performance of the CTE metric for predicting the model transferability ranking after the application of two unsupervised domain adaptation (UDA) approaches, Mean Teacher [65] (Tab. 23) and adaptive batch normalisation (AdaBN) [35] (Tab. 24). The tables show a sweep over a range of perturbation strengths and report the correlation scores per target dataset. In the main paper due to space limitations we reported the correlation scores averaged over the target datasets for a single perturbation strength. We investigated applying additive Gaussian noise to the input of the models with strengths varying in the range .
@@ 11.5 Post-UDA Perturbation Sweep
11.6 Final-Finetuned Correlation Transfer Metrics
For the supervised baseline methods compared against [47, 7, 79, 52, 27, 36, 74, 80] the goal of the transferability metrics is to predict the final supervised finetuned transfer performance of a model based on its initial pre-trained state. In the main paper we investigate the real world scenario of unsupervised model transfer, hence we are limited to assessing direct zero-shot application of pre-trained models to target data or assessing models after application of UDA approaches. We showed already in Tab. 1 that, even with supervision, existing transfer metrics fail at direct zero-shot transferability estimation in the semantic setting. How about in the setting of predicting the final supervised finetuned performance of models?
To investigate this we further trained a set of source models on the supervised training set of each of the mitochondria datasets in turn and then used the available ground truth to evaluate the performance of the fine-tuned models on the test set. Then using the pre-finetuning weights of the models we calculated all the base line transferability metrics and measured the correlation scores between pre-finetuned transferability metric and post-finetuned performance score. Tab. 25 shows the correlation scores split by target dataset, we can see that all the metrics fail to meaningfully correlate with final-finetuned performance. Further highlighting that existing transferability metrics cannot be reliably used for predicting the transfer of semantic segmentation models applied to biomedical data.
@@ 11.6 Final-Finetuned Correlation Transfer Metrics
| LogME | K | -0.13 | (0.59) | -0.02 | (0.98) | 0.35 | (0.17) | 0.00 | (1.00) |
| S | 0.05 | (0.84) | -0.01 | (0.97) | 0.41 | (0.22) | 0.13 | (0.66) | |
| P | 0.16 | (0.64) | -0.08 | (0.82) | 0.03 | (0.94) | 0.29 | (0.36) | |
| Hscore | K | -0.02 | (1.00) | 0.27 | (0.26) | 0.45 | (0.06) | 0.15 | (0.49) |
| S | 0.09 | (0.81) | 0.42 | (0.24) | 0.53 | (0.12) | 0.28 | (0.35) | |
| P | 0.20 | (0.56) | 0.21 | (0.54) | 0.41 | (0.20) | 0.17 | (0.60) | |
| GBC | K | 0.31 | (0.21) | 0.20 | (0.47) | 0.45 | (0.07) | 0.24 | (0.34) |
| S | 0.47 | (0.14) | 0.25 | (0.45) | 0.48 | (0.16) | 0.28 | (0.36) | |
| P | 0.24 | (0.47) | 0.53 | (0.10) | 0.41 | (0.22) | 0.45 | (0.14) | |
| RegHscore | K | -0.02 | (1.00) | 0.24 | (0.37) | 0.45 | (0.06) | 0.15 | (0.58) |
| S | 0.09 | (0.78) | 0.40 | (0.22) | 0.53 | (0.09) | 0.28 | (0.40) | |
| P | 0.20 | (0.56) | 0.21 | (0.54) | 0.41 | (0.21) | 0.17 | (0.60) | |
| NLEEP | K | 0.09 | (0.78) | -0.20 | (0.41) | 0.38 | (0.09) | 0.03 | (0.92) |
| S | 0.15 | (0.66) | -0.21 | (0.52) | 0.45 | (0.17) | -0.03 | (0.92) | |
| P | 0.11 | (0.75) | 0.10 | (0.76) | 0.45 | (0.17) | -0.02 | (0.95) | |
12 Instance Segmentation
In the following section we investigate ranking of instance segmentation models, each of the tables Tabs. 26, 27, 28, 29, 31, 30, 32 and 33 shows the correlation scores achieved by the CTE-ARS metric for sets of models applied directly to a single target dataset, across a range of perturbation strengths. In the main paper, due to space constraints, we show the correlation scores for a single perturbation strength averaged over the set of target datasets. We investigate the dependency of CTE on the strength of perturbation applied for both input and feature space perturbations. The tables show that CTE’s performance is stable across a wide range of perturbations.
12.1 Cells Perturbation Sweep
For cell instance segmentation we investigated applying additive Gaussian noise as an input perturbation between the strengths . We show the results for CTE-ARS in Tab. 26.
For feature space perturbation we investigated applying spatial DropOut equally to all layers of a network, where the strength of the perturbation is controlled by the proportion of feature maps dropped at each layer . We investigated DropOut proportions in the range . We show results for for both CTE-ARS in Tab. 27.
@@ 12.1 Cells Perturbation Sweep
12.2 Nuclei Perturbation Sweep
For nuclei instance segmentation we investigated applying a range of input augmentations; additive Gaussian noise, Gamma correction, Brightness adjustment and Contrast adjustment (as defined in Sec. 10). We investigated additive Gaussian noise with strengths in the range , Gamma correction with strengths in the range , where equals no adjustment, Brightness adjustment with strengths in the range and Contrast adjustment with strengths in the range , where equals no adjustment. We show the results for both CTE-ARS (Tabs. 28, 29, 30 and 31)
For feature space perturbation we investigated applying spatial DropOut only to the bottle neck layer of networks, where the strength of the perturbation is controlled by the proportion of feature maps dropped at each layer . We investigated DropOut proportions in the range . We show results for for CTE-ARS in Tab. 32.
We again observe stable correlation performance across a wide range of perturbation strengths. Although, as discussed in the main paper, Sec. 3.4, care should be taken to ensure perturbations remain ‘tolerable’ in the extreme perturbation case, too strong of a perturbation can lead to a loss of correlation between consistency based metrics and transfer performance. For example, the last row of Tab. 32 shows the strongest level of DropOut () applied can cause the CTE-ARS metric to become unstable on S_BIAD634.
| Metric | BBBC039 | Hoechst | S_BIAD895 | S_BIAD634 | |||||
@@ 12.2 Nuclei Perturbation Sweep
| Metric | BBBC039 | Hoechst | S_BIAD895 | S_BIAD634 | |||||
@@
| CTE-ARS Gamma ( 0.8 - 0.9) | K | 0.40 | (0.42) | 0.80 | (0.08) | -0.33 | (0.75) | 0.40 | (0.49) |
@@
| CTE-ARS Ctr ( 0.8 - 0.9) | K | 0.80 | (0.09) | 1.00 | (*) | 0.00 | (1.00) | 0.80 | (0.08) |
| S | 0.90 | (0.09) | 1.00 | (*) | -0.20 | (0.92) | 0.90 | (0.07) | |
| P | 0.85 | (0.07) | 0.94 | (*) | 0.18 | (0.82) | 0.83 | (0.08) | |
| CTE-ARS Ctr ( 0.9 - 0.95) | K | 0.80 | (0.09) | 1.00 | (*) | 0.67 | (0.33) | 0.80 | (0.10) |
| S | 0.90 | (0.09) | 1.00 | (*) | 0.80 | (0.33) | 0.90 | (0.09) | |
| P | 0.96 | (*) | 0.99 | (**) | 0.52 | (0.48) | 0.87 | (0.05) | |
| CTE-ARS Ctr ( 1.05 - 1.1) | K | 0.74 | (0.13) | 1.00 | (*) | 0.67 | (0.33) | 0.60 | (0.25) |
| S | 0.82 | (0.13) | 1.00 | (*) | 0.80 | (0.33) | 0.80 | (0.13) | |
| P | 0.99 | (**) | 0.95 | (*) | 0.58 | (0.42) | 0.80 | (0.10) | |
| CTE-ARS Ctr ( 1.1 - 1.2) | K | 0.80 | (0.11) | 0.60 | (0.23) | 0.33 | (0.75) | 0.80 | (0.07) |
| S | 0.90 | (0.10) | 0.70 | (0.20) | 0.40 | (0.75) | 0.90 | (0.09) | |
| P | 0.83 | (0.08) | 0.81 | (0.10) | 0.38 | (0.62) | 0.90 | (*) | |
| Metric | BBBC039 | Hoechst | S_BIAD895 | S_BIAD634 | |||||
| pval. | pval. | pval. | pval. | ||||||
| CTE-ARS Brt ( 0.0 - 0.05) | K | 0.60 | (0.22) | 0.60 | (0.26) | 0.00 | (1.00) | 0.80 | (0.10) |
| S | 0.70 | (0.25) | 0.70 | (0.21) | 0.20 | (0.92) | 0.90 | (0.09) | |
| P | 0.80 | (0.10) | 0.83 | (0.08) | 0.77 | (0.23) | 0.72 | (0.17) | |
| CTE-ARS Brt ( 0.05 - 0.1) | K | 0.40 | (0.46) | 0.80 | (0.11) | 0.33 | (0.75) | 0.80 | (0.07) |
| S | 0.60 | (0.32) | 0.90 | (0.10) | 0.40 | (0.75) | 0.90 | (0.10) | |
| P | 0.69 | (0.19) | 0.92 | (*) | 0.80 | (0.20) | 0.84 | (0.08) | |
| CTE-ARS Brt ( 0.1 - 0.2) | K | 0.60 | (0.23) | 0.80 | (0.09) | 0.67 | (0.33) | 0.40 | (0.49) |
| S | 0.70 | (0.23) | 0.90 | (0.10) | 0.80 | (0.33) | 0.50 | (0.40) | |
| P | 0.81 | (0.10) | 0.85 | (0.07) | 0.99 | (*) | 0.43 | (0.47) | |
| Metric | BBBC039 | Hoechst | S_BIAD895 | S_BIAD634 | |||||
| pval. | pval. | pval. | pval. | ||||||
| CTE-ARS DropOut ( 0.05) | K | 0.40 | (0.43) | 0.80 | (0.08) | 1.00 | (0.08) | 0.80 | (0.12) |
@@
12.3 Covid_IF External Models Perturbation Sweep
In the main paper in Fig. 4 we also ranked the transferability of a set of publicly available models [51, 3, 53] for instance cell segmentation on the Covid_IF [53] dataset. The models all have very different instantiation processes, but through calculating the transferability metric in output segmentation space we are able to directly compare the varied set of models. The Micro-SAM models [3] perform Automatic Instance Segmentation (AIS), which utilises an additionally trained decoder to predict a three channel output: the distance to the object centre, the distance to the object boundary and foreground probabilities. The predictions are then instantiated using seeded watershed. Cellpose-SAM [51] use a custom decoder to directly predict vector flow fields, which can then be converted to instance segmentations. ‘Powerful-chipmunk’ is a specialist U-Net model taken from the Bioimage model zoo [49], the model is trained specifically on the Covid_IF data and predicts object foreground and boundary predictions, which are then converted to instances using watershed and GASP.
In Fig. 4 from the main paper we show the correlation between CTE-ARS and mSA performance for a single additive Gaussian noise perturbation strength. In Tab. 33 we investigate a range of Gaussian noise strengths .

### v3

@@ 0.E.3 Baseline Transferability Metrics
| NuNo | K | 0.49 | (*) | 0.2 | (0.28) | 0.06 | (0.85) | -0.06 | (0.87) |
| S | 0.68 | (*) | 0.20 | (0.51) | -0.03 | (0.92) | -0.04 | (0.90) | |
| P | 0.52 | (*) | 0.10 | (0.72) | -0.05 | (0.86) | -0.04 | (0.90) | |
| Dispersion | K | 0.03 | (0.94) | -0.07 | (0.78) | -0.06 | (0.73) | 0.00 | (1.00) |
| S | 0.05 | (0.87) | -0.13 | (0.66) | -0.12 | (0.66) | -0.08 | (0.81) | |
| P | -0.07 | (0.81) | -0.25 | (0.36) | -0.20 | (0.49) | -0.17 | (0.60) | |
@@
| NuNo | K | 0.33 | (0.36) | -0.43 | (0.23) | -0.14 | (0.82) | 0.60 | (0.13) |
| S | 0.57 | (0.16) | -0.43 | (0.37) | -0.11 | (0.85) | 0.66 | (0.19) | |
| P | 0.15 | (0.75) | -0.48 | (0.28) | -0.04 | (0.93) | 0.68 | (0.14) | |
| Dispersion | K | 0.24 | (0.54) | -0.24 | (0.56) | -0.24 | (0.53) | -0.47 | (0.29) |
| S | 0.25 | (0.58) | -0.43 | (0.36) | -0.39 | (0.44) | -0.60 | (0.26) | |
| P | 0.43 | (0.33) | -0.02 | (0.97) | -0.62 | (0.14) | -0.56 | (0.25) | |
0.E.4 Multiclass Toothfairy2 Perturbation Sweep
For multiclass semantic segmentation we built a diverse set of models with . All but one of the models were trained on the Toothfairy2 datasets taken from the MICCAI2024 challenge [2024TMI, 2025CVPR], the exception being the generalist TotalSegmentator [wasserthal_totalsegmentator_2023] model. The model set comprised of a wide range of architectures: 2D nnU-Net, residual-encoder 3D nnU-Net, SwinUNETR, TotalSegmentator, UMamba and VMamba.
We investigated applying both additive Gaussian noise (controlled by ) and Gamma Correction (controlled by ) as input perturbations. For feature space perturbation we investigated applying spatial DropOut to the bottleneck and skip connection layers of the network, where the strength of the perturbation is controlled by the proportion of feature maps dropped at each layer . We report the performance of both CMR-EI (Tab.˜14) and CMR-NHD (Tab.˜15).
@@ 0.E.4 Multiclass Toothfairy2 Perturbation Sweep
| CMR-EI Gauss ( 2.5x) | K | 0.93 | (**) | 0.86 | (*) | 0.65 | (*) | 0.40 | (0.2) | 0.71 | (*) |
| S | 0.97 | (**) | 0.85 | (0.1) | 0.85 | (**) | 0.44 | (0.1) | 0.84 | (*) | |
| P | 0.96 | (*) | 0.94 | (**) | 0.73 | (*) | 0.39 | (0.3) | 0.90 | (*) | |
| CMR-EI Gauss ( 3.0x) | K | 0.90 | (**) | 0.84 | (**) | 0.65 | (*) | 0.43 | (0.2) | 0.71 | (*) |
| S | 0.97 | (**) | 0.89 | (**) | 0.85 | (**) | 0.44 | (0.2) | 0.86 | (*) | |
| P | 0.96 | (*) | 0.92 | (**) | 0.59 | (0.2) | 0.42 | (0.2) | 0.90 | (*) | |
| CMR-EI Gamma ( 0.2) | K | 0.93 | (**) | 0.77 | (*) | 0.65 | (0.1) | 0.57 | (0.1) | 0.78 | (*) |
| S | 0.96 | (**) | 0.88 | (**) | 0.74 | (0.2) | 0.61 | (0.1) | 0.88 | (0.1) | |
| P | 0.95 | (*) | 0.71 | (*) | 0.66 | (0.1) | 0.50 | (0.1) | 0.70 | (0.2) | |
| CMR-EI Gamma ( 0.5) | K | 0.97 | (**) | 0.77 | (**) | 0.68 | (0.1) | 0.48 | (0.2) | 0.81 | (**) |
| S | 0.98 | (**) | 0.90 | (**) | 0.80 | (*) | 0.51 | (0.2) | 0.89 | (**) | |
| P | 0.94 | (*) | 0.83 | (**) | 0.75 | (*) | 0.42 | (0.3) | 0.89 | (**) | |
| CMR-EI Gamma ( 2.0) | K | 1.00 | (**) | 0.71 | (*) | 0.43 | (0.3) | 0.43 | (0.2) | 0.81 | (**) |
@@
0.E.5 Post-UDA Perturbation Sweep
In this section we investigate the performance of the CMR metric for predicting the model ttarget performance ranking after the application of two unsupervised domain adaptation (UDA) approaches, Mean Teacher [tarvainen_mean_2018] (Tab.˜16(a)) and adaptive batch normalisation (AdaBN) [li_adaptive_2018] (Tab.˜16(b)). The tables show a sweep over a range of perturbation strengths and report the correlation scores per target dataset. In the main paper due to space limitations we reported the correlation scores averaged over the target datasets for a single perturbation strength. We investigated applying additive Gaussian noise to the input of the models with strengths varying in the range .
@@ 0.E.5 Post-UDA Perturbation Sweep
Tab.˜16(b) shows that the correlation score of our CMR approach remains very stable over a wide range of perturbation strengths. As noted in the main paper CMR ranking strongly correlates to F1-score ranking across three out of the four Mitochondria datasets, but struggles with the VNC dataset (but still beats the Transfer Score baseline see Tab.˜4). We wanted to investigate this apparent underperformance on VNC further.
Firstly, it is important to note that VNC is by far the smallest dataset thus reducing its statistical robustness. The entire VNC dataset only 20 slices of pixels. As noted in Sec.˜0.A.1, due to the small dataset size in direct zero-shot application we purposely disregarded the source-to-source VNC model transfer and instead used the full 20 slices as test data for models trained on other sources. However, in post-UDA ranking analysis this is no longer possible as we must split the data into train/test splits so that models can be adapted on the target training set and then tested on the test set. Hence, for the post-UDA ranking the models transferred to VNC are assessed on only 2 slices.
However, when we look into the errors made by the CMR ranking for post-UDA transfer to the VNC dataset we see that models trained on the EPFL dataset seem to be outliers with unexpectedly high consistency scores despite low performance. When investigating this further it seemed there was class confusion, see Fig.˜5, in the predictions of the EPFL models, where ‘none mitochondria’ objects are consistently but erroneously segmented by the transferred EPFL source models. This results in low performance F1-score, but high consistency score.
As noted in the main paper consistency regularisation in Mean Teacher training likely reinforces this class confusion between visually similar, but distinct structures present in EPFL and VNC, violating the assumption that source and target models address the same task. Hence, it is not surprising that CMR ranking after AdaBN UDA, which lacks a consistency regularisation component, performs better on the VNC target dataset, as shown in Tab.˜16(b).
0.E.6 Transferability Metrics: Final-Finetuned Correlation
For the supervised baseline methods compared against [nguyen_leep_2020, bao_information-theoretic_2022, yang_pick_2023, pandy_transferability_2021, ibrahim_newer_2023, li_ranking_2020, wang_how_2023, you_logme_2021] the goal of the transferability metrics is to predict the final supervised finetuned transfer performance of a model based on its initial pre-trained state. In the main paper we investigate the real world scenario of unsupervised model transfer, hence we are limited to assessing direct zero-shot application of pre-trained models to target data or assessing models after application of UDA approaches. We showed already in Tab.˜1 that, even with supervision, existing transfer metrics fail at direct zero-shot target performance ranking in the semantic setting. How about in the setting of predicting the final supervised finetuned performance of models?
To investigate this we further trained a set of source models on the supervised training set of each of the mitochondria datasets in turn and then used the available ground truth to evaluate the performance of the fine-tuned models on the test set. Then using the pre-finetuning weights of the models we calculated all the base line transferability metrics and measured the correlation scores between pre-finetuned transferability metric and post-finetuned performance score. Tab.˜17 shows the correlation scores split by target dataset, we can see that all the metrics fail to meaningfully correlate with final-finetuned performance. Further highlighting that existing transferability metrics cannot be reliably used for predicting the transfer of semantic segmentation models applied to biomedical data.
@@ 0.E.6 Transferability Metrics: Final-Finetuned Correlation
Appendix 0.F Instance Segmentation
In the following section we investigate ranking of instance segmentation models, each of the tables Tabs.˜18, 19, 20(a), 20(d), 20(b), 20(c), 21 and 22 shows the correlation scores between our CMR-ARS metric and instance segmentation performance metrics for sets of models applied directly to a single target dataset, across a range of perturbation strengths. In the main paper, due to space constraints, we show the correlation scores for a single perturbation strength averaged over the set of target datasets. We investigate the dependency of CMR on the strength of perturbation applied for both input and feature space perturbations. The tables show that CMR’s performance is stable across a wide range of perturbations.
0.F.1 Cells Perturbation Sweep
For cell instance segmentation we investigated ranking a diverse set of domain specialist models with . The models were trained on three different source datasets FlyWing [leal-taixe_benchmark_2019], Ovules [wolny_accurate_2020] and PNAS [willis_cell_2016] and include 2D U-Nets (trained both with and without augmentations) and Residual 2D U-Nets.
We investigated applying additive Gaussian noise as an input perturbation between the strengths . We show the results for CMR-ARS in Tab.˜18.
For feature space perturbation we investigated applying spatial Test-time DropOut (TTD) equally to all layers of a network, where the strength of the perturbation is controlled by the proportion of feature maps dropped at each layer . We investigated DropOut proportions in the range . We show results for for both CMR-ARS in Tab.˜19.
@@ 0.F.1 Cells Perturbation Sweep
0.F.2 Nuclei Perturbation Sweep
For nuclei instance segmentation we trained a set of domain specialist models with . The models all have a 2D U-Net architecture and were trained on five datasets; BBBC039 [ljosa_annotated_2012], HeLaCytoNuc [de_helacytonuc_2024], Hoechst [arvidsson_annotated_2023], S_BIAD895/SB-895 [von_chamier_democratising_2021] and SBIAD1410 [hawkins_rescu-nets_2024]. The source models were then transferred to the test sets of four target datasets BBBC039 [ljosa_annotated_2012], Hoechst [arvidsson_annotated_2023], S_BIAD895 [von_chamier_democratising_2021] and S_BIAD634/SB-634 [kromp_annotated_2020].
We investigated applying a range of input augmentations; additive Gaussian noise, Gamma correction, Brightness adjustment and Contrast adjustment (as defined in Appendix˜0.C). We investigated additive Gaussian noise with strengths in the range , Gamma correction with strengths in the range , where equals no adjustment, Brightness adjustment with strengths in the range and Contrast adjustment with strengths in the range , where equals no adjustment. We show the results for both CMR-ARS (Tabs.˜20(a), 20(d), 20(c) and 20(b))
For feature space perturbation we investigated applying spatial DropOut only to the bottle neck layer of networks, where the strength of the perturbation is controlled by the proportion of feature maps dropped at each layer . We investigated DropOut proportions in the range . We show results for for CMR-ARS in Tab.˜21.
We again observe stable correlation performance across a wide range of perturbation strengths. Although, as discussed in the main paper, Sec.˜3.3, care should be taken to ensure perturbations remain ‘tolerable’ in the extreme perturbation case, too strong of a perturbation can lead to a loss of correlation between consistency based metrics and transfer performance. For example, the last row of Tab.˜21 shows the strongest level of DropOut () applied can cause the CMR-ARS metric to become unstable on S_BIAD634.
| Metric | BBBC039 | Hoechst | SB-895 | SB-634 | |||||
@@ 0.F.2 Nuclei Perturbation Sweep
| Metric | BBBC039 | Hoechst | SB-895 | SB-634 | |||||
@@
| CMR-ARS Brt, | K | 0.60 | (0.22) | 0.60 | (0.26) | 0.00 | (1.00) | 0.80 | (0.10) |
| S | 0.70 | (0.25) | 0.70 | (0.21) | 0.20 | (0.92) | 0.90 | (0.09) | |
| P | 0.80 | (0.10) | 0.83 | (0.08) | 0.77 | (0.23) | 0.72 | (0.17) | |
| CMR-ARS Brt, | K | 0.40 | (0.46) | 0.80 | (0.11) | 0.33 | (0.75) | 0.80 | (0.07) |
| S | 0.60 | (0.32) | 0.90 | (0.10) | 0.40 | (0.75) | 0.90 | (0.10) | |
| P | 0.69 | (0.19) | 0.92 | (*) | 0.80 | (0.20) | 0.84 | (0.08) | |
| CMR-ARS Brt, | K | 0.60 | (0.23) | 0.80 | (0.09) | 0.67 | (0.33) | 0.40 | (0.49) |
| S | 0.70 | (0.23) | 0.90 | (0.10) | 0.80 | (0.33) | 0.50 | (0.40) | |
| P | 0.81 | (0.10) | 0.85 | (0.07) | 0.99 | (*) | 0.43 | (0.47) | |
| Metric | BBBC039 | Hoechst | SB-895 | SB-634 | |||||
| pval. | pval. | pval. | pval. | ||||||
| CMR-ARS Ctr, | K | 0.80 | (0.09) | 1.00 | (*) | 0.00 | (1.00) | 0.80 | (0.08) |
| S | 0.90 | (0.09) | 1.00 | (*) | -0.20 | (0.92) | 0.90 | (0.07) | |
| P | 0.85 | (0.07) | 0.94 | (*) | 0.18 | (0.82) | 0.83 | (0.08) | |
| CMR-ARS Ctr, | K | 0.80 | (0.09) | 1.00 | (*) | 0.67 | (0.33) | 0.80 | (0.10) |
| S | 0.90 | (0.09) | 1.00 | (*) | 0.80 | (0.33) | 0.90 | (0.09) | |
| P | 0.96 | (*) | 0.99 | (**) | 0.52 | (0.48) | 0.87 | (0.05) | |
| CMR-ARS Ctr, | K | 0.74 | (0.13) | 1.00 | (*) | 0.67 | (0.33) | 0.60 | (0.25) |
| S | 0.82 | (0.13) | 1.00 | (*) | 0.80 | (0.33) | 0.80 | (0.13) | |
| P | 0.99 | (**) | 0.95 | (*) | 0.58 | (0.42) | 0.80 | (0.10) | |
| CMR-ARS Ctr, | K | 0.80 | (0.11) | 0.60 | (0.23) | 0.33 | (0.75) | 0.80 | (0.07) |
| S | 0.90 | (0.10) | 0.70 | (0.20) | 0.40 | (0.75) | 0.90 | (0.09) | |
| P | 0.83 | (0.08) | 0.81 | (0.10) | 0.38 | (0.62) | 0.90 | (*) | |
| Metric | BBBC039 | Hoechst | SB-895 | SB-634 | |||||
| pval. | pval. | pval. | pval. | ||||||
| CMR-ARS Gamma, | K | 0.40 | (0.42) | 0.80 | (0.08) | -0.33 | (0.75) | 0.40 | (0.49) |
@@
| CMR-ARS DropOut ( 0.05) | K | 0.40 | (0.43) | 0.80 | (0.08) | 1.00 | (0.08) | 0.80 | (0.12) |
@@
0.F.3 Covid_IF External Models Perturbation Sweep
In the main paper in Fig.˜4 we also ranked the predicted target performance of a set of publicly available models [pachitariu_cellpose-sam_2025, archit_segment_2025, pape_microscopy-based_2021] for instance cell segmentation on the Covid_IF [pape_microscopy-based_2021] dataset. The models all have very different instantiation processes, but through calculating the ranking metric in output segmentation space we are able to directly compare the varied set of models. The Micro-SAM models [archit_segment_2025] perform Automatic Instance Segmentation (AIS), which utilises an additionally trained decoder to predict a three channel output: the distance to the object centre, the distance to the object boundary and foreground probabilities. The predictions are then instantiated using seeded watershed. Cellpose-SAM [pachitariu_cellpose-sam_2025] use a custom decoder to directly predict vector flow fields, which can then be converted to instance segmentations. ‘Powerful-chipmunk’ is a specialist U-Net model taken from the Bioimage model zoo [ouyang_bioimage_2022], the model is trained specifically on the Covid_IF data and predicts object foreground and boundary predictions, which are then converted to instances using watershed and GASP.
In Fig.˜4 from the main paper we show the correlation between CMR-ARS and mSA performance for a single additive Gaussian noise perturbation strength. In Tab.˜22 we investigate a range of Gaussian noise strengths .